OpenStudy (anonymous):

y=3x^2-4x-1 write in standard form

OpenStudy (coconutjj):

3x^2-4x=1

OpenStudy (freckles):

y=ax^2+bx+c is standard form for a parabola do you mean vertex form?

OpenStudy (anonymous):

Ax^2 + Bx + C = 0 is standard form I think